TAC Face-to-Face Meeting. DanaBalser briefly summarized the TAC postmortem. The main issues included (1) High SRP loads, (2) Competing proposals, (3) Large proposals and data management, (4) HSA long term projects, and (5) Inverted SRP scores. We spend most of the time discussing (1). The TAC suggested creating two panels in place of ETP. There were questions of a plan B (e.g., if ETP shrinks or AGN expands). Most think it likely that ETP will not significantly shrink (we need at least 20 per SRP) or that AGN will expand too much since it is VLA configuration dependent. BarryClark recalled his earlier suggestion of allowing the proposers to select keywords that would then be used to generate categories. This could be a longer term solution, maybe with the new tool suite, but has the disadvantage of potentially more work for observatory staff on the backside. RickLively will try to have something ready for testing soon so we can assess if this will be straight forward or not. GarethHunt noted that changing the SRP categories will create some software work with metrics.
Directors' Review - Tuesday, 6 November 2018, 11:00-14:00 EST (09:00-12:00 MST).
Logistics. The time is Tuesday, 6 November 2018, 11:00-14:00 EST (09:00-12:00 MST). The plan is to use the Director office in CV and GB and a reserved room in SO. ToneyMinter mentioned that KarenONeil will not be available until 1pm EST. We will shift the GBT until last.
GBT Info. Already sent. It was noted that the PI name of one of the Large proposals was not correct. It was not clear why this happened.
VLA Info. AmyMioduszewski reported that she should be done by COB today. The plan is to converge with BarryClark and create a master version that includes the ANY configuration mapping. This should be okay since the disposition letters do not indicate the semester. BryanButler noted that they have not had the RSRO meeting so this will need to be included later. Since the RSRO meeting is tomorrow the outcomes will be available before the Directors' review.
VLBA Info. MarkClaussen reported that he should be completed by the COB today.
Metrics. AnandCrossley reported that he should be completed by the COB today.
PST Tickets Status. RickLively reported that he is a little behind schedule. He will try to have the major items soon. He plans to work on the two SRPs item first since this has implications for the CfP and recruitment. LorantSjouwerman and BryanButler converged on some smaller VLA resource items.
SRP Recruitment. MarkClaussen reported no action here, but LewisBall reported to him some rejections on TAC recruitment. We briefly discussed the implications of adding an SRP. Currently there are 7 SRP + 1 SRP chair for ETP. Normally we would need to recruit one more reviewer for a total of 5 normal + 3 supplemental. If we go to two SRPs then we would need (5+1)*2 = 12 reviewers and thus need to recruit 4 people.